Solve the following system of equations.
4a+7b=-24
8a+3c=2
6b+3c=-30

Answer by TimothyLamb(4379)

4a + 7b = -24
8a + 3c = 2
6b + 3c = -30
---
put the system of linear equations into standard form
---
4x + 7y = -24
8x + 3z = 2
6y + 3z = -30
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
x= 1
y= -4
z= -2
---
answer:
a= 1
b= -4
c= -2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php